# WeChat-YATT: A Simple, Scalable and Balanced RLHF Trainer

### v1

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has emerged as a prominent paradigm for training large language models and multimodal systems. Despite notable advances enabled by existing RLHF training frameworks, significant challenges remain in scaling to complex multimodal workflows and adapting to dynamic workloads. In particular, current systems often encounter limitations related to controller scalability when managing large models, as well as inefficiencies in orchestrating intricate RLHF pipelines—especially in scenarios that require dynamic sampling and resource allocation. In this paper, we introduce WeChat-YATT (Yet Another Transformer Trainer in WeChat), a simple, scalable, and balanced RLHF training framework specifically designed to address these challenges. WeChat-YATT features a parallel controller programming model that enables flexible and efficient orchestration of complex RLHF workflows, effectively mitigating the bottlenecks associated with centralized controller architectures and facilitating scalability in large-scale data scenarios. In addition, we propose a dynamic placement schema that adaptively partitions computational resources and schedules workloads, thereby significantly reducing hardware idle time and improving GPU utilization under variable training conditions. We evaluate WeChat-YATT across a range of experimental scenarios, demonstrating that it achieves substantial improvements in throughput compared to state-of-the-art RLHF training frameworks. Furthermore, WeChat-YATT has been successfully deployed to train models supporting WeChat product features for a large-scale user base, underscoring its effectiveness and robustness in real-world applications.
@@ 1 Introduction
Transformer-based large models and their multimodal extensions have achieved remarkable progress, demonstrating state-of-the-art performance across a wide range of natural language processing, understanding, and generation tasks [1, 2, 3, 4, 5, 6, 7, 8, 9]. These models, characterized by their massive parameter counts and ability to process extremely long context sequences, have become the foundation for many advanced AI applications. However, the unprecedented scale of LLMs has introduced significant challenges for efficient training, traditional data-parallel training strategies are often insufficient to handle the memory and computational demands of such large models. To address these challenges, the research community has developed a variety of hybrid parallelism techniques [10, 11, 12, 13, 14]. These approaches enable the training of models that would otherwise exceed the memory and compute capabilities of individual devices, by partitioning both the model and the data across multiple GPUs or nodes.
@@
The integration of Reinforcement Learning from Human Feedback (RLHF) has significantly transformed the training paradigms of large language models [15, 16, 17]. By embedding human preferences into the optimization process, RLHF necessitates the coordination of multiple models—such as reward models, policy models, and reference models—within a unified training pipeline [18, 19, 20]. This multi-model architecture substantially amplifies the number of parameters and the volume of intermediate activations that must be handled, thereby straining the capabilities of current distributed training infrastructures. The scalability challenges are further exacerbated in multimodal RLHF scenarios. The incorporation of diverse data modalities (e.g., text, images, audio) not only increases the heterogeneity and volume of training data but also intensifies the demands on inter-node communication and memory bandwidth. As a result, the system faces pronounced communication overheads and memory bottlenecks, which impede efficient scaling to larger datasets and model sizes. Addressing these bottlenecks is critical for advancing the practical deployment of RLHF in increasingly complex and data-rich environments.
@@
Sample efficiency is a critical factor in the RLHF training process, as it directly influences convergence rates and overall model performance [21]. While flexible sampling strategies have been shown to accelerate convergence, most existing RLHF training frameworks lack robust support for efficient and adaptive sampling mechanisms. Although hybrid parallelism theoretically enables substantial scalability, practical implementations often encounter significant communication overhead as the degree of parallelism increases. Specifically, as the number of models and the length of context sequences expand, the costs associated with synchronizing parameters and activations across distributed devices can escalate rapidly, thereby diminishing the anticipated benefits of parallelization. To address these challenges, recent work such as [22] has conceptualized the RLHF training workflow as a control flow problem, employing techniques such as multi-cluster, multi-model partial colocated, and memory swapping to alleviate peak GPU memory consumption. Similarly, other systems [23, 24] have explored alternative strategies, with [24] further enhancing throughput by integrating both single- and multi-controller architectures alongside statically optimized placement schemes. These advancements underscore the ongoing need for innovative approaches to improve sampling efficiency and resource utilization in large-scale RLHF training.
@@
We have open-sourced WeChat-YATT at https://www.github.com/tencent/WeChat-YATT. We believe that this framework advances the state-of-the-art in RLHF training by effectively addressing the critical bottlenecks of controller scalability and resource placement. We hope that WeChat-YATT will serve as a robust foundation for both academic research and industrial applications, fostering further innovation and collaboration within the community.
@@ 2.1 Distributed Training of Deep Models
Data Parallelism involves splitting the training data across multiple devices (such as GPUs), where each device holds a copy of the model and processes a different mini-batch of data. Gradients are averaged (or summed) across devices to update the model parameters synchronously [25, 26]. Tensor parallelism splits individual layers or tensors of the model across multiple devices. Each device is responsible for computing a portion of the operations within a layer, enabling the training of very large models that cannot fit into a single device’s memory [10]. Pipeline parallelism divides the model into sequential stages, with each stage placed on a different device. Mini-batches are split into micro-batches and passed through the pipeline, allowing different devices to work on different parts of the input simultaneously, improving hardware utilization [11, 27, 28].
ZeRO is an optimization technique that partitions model states (such as optimizer states, gradients, and parameters) across devices to minimize memory redundancy. This enables the training of much larger models by reducing memory consumption per device [29, 30].
@@
Expert Parallelism is used in Mixture-of-Experts (MoE) models, where different subsets of the model (experts) are distributed across devices. During training and inference, only a subset of experts is activated for each input, enabling efficient scaling and utilization of resources [13, 33].
@@ 2.2 RLHF Workflow
Most RLHF workflows largely follow this 4-stage workflow [18, 19, 40, 16, 15, 41, 20](Sometimes, the first three stages are referred to as "rollout" (Figure 1(a)):
@@
    Stage 3(Preparation): Using generated responses and rewards, the critic computes their values, the policy and reference policy computes their reference log probabilities.
@@
To simplify the implementation, most systems use a single controller to manage data transfer, intermediate results, and workflow orchestration through a central point [23, 24]. However, this centralized approach can sometimes become a bottleneck—especially when the trainer itself has high memory requirements.
@@ 2.3 RLHF Placement
  • Co-locate: Multiple models are placed on the same set of GPUs, sharing GPU memory. To avoid out-of-memory (OOM) errors, which can easily occur if large models run simultaneously, these colocated models are executed one after another in a time-sharing fashion.
Although the [24] proposed methods for calculating placement based on workload, both the open-source implementations of [23] and [24] adopt a co-locate all models strategy. The partial colocated approach introduces overhead from swapping among multiple models. For example, swapping a 32B model could take nearly a minute, and updating weights may take tens of seconds. In typical GRPO training, this overhead is not a major issue, since the primary bottlenecks in RLHF training are the generation and the forward/backward passes of modeling. Compared to the tens of minutes spent on rollout and training, model swapping is not the system bottleneck.
Some works [44, 45] have adopted a co-existing approach, which effectively avoids the overhead of model swapping. However, to mask device idle time during workflow state transitions, these methods introduce varying degrees of asynchronous overlap between the rollout and training stages. This can lead to suboptimal model convergence; for example, [44] introduces staleness-aware training to address this issue. In our experience, many algorithm engineers find that this tradeoff adds extra cognitive burden.
@@ 3.1 Overview
Figure 2 presents a high-level overview of the proposed WeChat-YATT framework. WeChat-YATT supports two distinct resource placement architectures: partial colocated and fully colocated.
In the partial colocated architecture, the Sampler and the GenRM (Generative Reward Model) components are deployed independently and interact asynchronously. The allocation of computational resources to these two components directly influences the throughput during the rollout phase, a placement method that will be examined in detail in Section 3.3. Within this architecture, the actor model leverages all available machine resources across the cluster, while parallel controllers correspond to the ranks of model-parallel worker heads associated with the Actor. The training loop is initiated by wake-up remote procedure calls (RPCs), which activate both the Sampler and GenRM. Subsequently, parallel controllers retrieve data from the distributed GDataset, thereby mitigating excessive memory consumption on any single machine. Notably, both the Sampler and GenRM operate asynchronously: upon completion of each microbatch, the Sampler immediately transmits the generated data to the GenRM without impeding ongoing generation processes. Once the rollout phase concludes, the Actor instructs other components to enter a sleep state, offloading their memory, and then loads the actor model to commence the training phase.
In contrast, the fully colocated architecture adopts a more integrated approach, wherein all components share the resources of all machines. In this configuration, the GenRM must wait until the Sampler has completed the generation of all samples. Following this, the Sampler offloads its parameters to disk, and the GenRM is loaded into memory to initiate the reward computation process. Upon completion of reward computation, the Actor is loaded into memory while the GenRM is offloaded. This workflow is inherently synchronous, with transitions between components occurring only after the preceding process has fully completed. Consequently, in scenarios characterized by frequent long-tail effects, the fully colocated architecture may not offer optimal performance.
@@ 3.3 Dynamic Placement
However, dynamic sampling scenarios introduce additional complexity. As described in [40], prompts with accuracy scores of 1 or 0 (i.e., batches with uniformly high or low rewards) are filtered out, triggering resampling. This process necessitates additional rounds of generation and reward computation, as illustrated in Figure 1(b).
@@
To address the challenges associated with model swapping overhead and the exacerbated long-tail effect, we have developed a dynamic placement strategy that integrates that dynamically determine resource allocation in the training process under patial colocated WeChat-YATT paradigm.
@@
When the workload is dynamic, static estimation methods such as those proposed in [24] become inadequate for determining optimal resource placement. It is inherently difficult to predict how the output length of the model will evolve over the course of training. Moreover, as the model’s reasoning responses become longer, it becomes increasingly challenging to anticipate the corresponding generation length required for reward computation. This unpredictability further complicates the design of efficient placement strategies in dynamic training environments.
@@ 4.4 Workload Balancing
As noted in [54, 55], workload imbalance can create significant bubbles during training, becoming a major bottleneck. When sequence lengths are long, the main training bottleneck is attention computation, whose complexity is $s^{2}$ for a sequence of length $s$, while the cost of packing is much lower. For example, if a sequence is formed by concatenating two sequences of length $\frac{s}{2}$, the total computation is $\frac{s^{2}}{2}$, leading to severe load imbalance and device underutilization. This issue is even more pronounced with post-training data, which often varies greatly in length.
Compared to complex combinatorial optimization approaches, we found a much simpler solution for workload balancing: instead of sequence packing, we simply sort data by simulated workload ($s^{2}$), greatly simplifying the process and even improving computational efficiency.
@@ 4.5 Distributed Attention
Currently, most implementations achieve context parallelism using ring attention for both causal and full attention. To support more complex attention masks (e.g., Gemma-3 [56]), inspired by Llama-3 [2], we implemented distributed attention based on CCL all-gather, where we first all-gather the key (K) and value (V) tensors, and then compute attention output for the local query (Q) tensor chunk.
@@ 5.1.1 Co-locate Generative Reward Model
In the full colocated WeChat-YATT setup, we conducted experiments on the generative reward model compared with VeRL under two distinct scenarios. The first scenario features a balanced parameter size between the generative reward model and the actor model, while the second scenario considers a highly unbalanced configuration. As VeRL currently supports only external calls to the generative reward model, we allocated computational resources in proportion to the parameter size ratio between the generative reward and actor models.
For the first comparation scenario setup, we selected the VeRL Generative Reward demo. In this configurationof VeRL, GenRM-CI-Test-1.5B serves as the external reward model, utilizing 8 GPUs for inference. SGlang is employed as the inference engine, with a data parallelism size (dp-size) of 8. An additional 8 GPUs are allocated to deploy the actor model, Qwen2.5-3B-Instruct, with a tensor parallelism size (tp-size) of 2. For the training backend, we adopt the state-of-the-art Megatron-LM framework. To ensure a fair comparison, the WeChat-YATT setup mirrors the VeRL configuration, employing a full colocated deployment with 16 GPUs.
@@
We further evaluated performance in a scenario involving a large generative reward model, Qwen2.5-Math-72B-Instruct, paired with a smaller actor model, Qwen2.5-Math-1.5B. This configuration corresponds to the generative reward model demonstration provided in our open-source codebase. For the VeRL setup, 24 GPUs were allocated for inference, while an additional 8 GPUs were dedicated to actor model rollout generation and training. In contrast, WeChat-YATT employed a full colocated configuration utilizing 32 GPUs, leveraging an offload/onload strategy. Megatron-LM was used as the training backend in both cases.
@@ 5.1.2 Dynamic Sampling
To assess the performance of the partial colocated WeChat-YATT framework in a dynamic sampling context, we conducted experiments in which varying sample rejection rates were applied during the rollout phase. Consistent with previous experimental settings, GenRM-CI-Test-1.5B was employed as the generative reward model, while Qwen2.5-3B-Instruct served as the actor model. The sample rejection rates are varied from 10% to 40%, and the framework need to re-sample rejected rollouts until all samples are accepted. The primary metric of interest was the time consumption observed during the rollout phase under these different rejection rates.
As shown in Table LABEL:tab:reject_rates, the rollout time for both frameworks increases as the sample rejection rate rises. However, the partial colocated WeChat-YATT consistently demonstrates lower rollout time across all rejection rates when compared to the fully colocated approach. Specifically, the partial colocated framework achieves a reduction in rollout time usage ranging from 16.2% to 24.1% relative to the fully colocated configuration.
This performance improvement can be attributed to the asynchronous interaction between the actor rollouts and the generative reward model in the partial colocated design. In this setup, each completed microbatch of actor rollouts is immediately dispatched to the standalone generative reward model in an asynchronous manner. This process effectively mitigates the long-tail latency that often constrains throughput in the fully colocated WeChat-YATT framework. Furthermore, the frequent context switching between actor rollouts and the generative model in the fully colocated setup introduces additional overhead, which is substantially reduced in the partial colocated architecture.

### v3

# WeChat-YATT: A Scalable, Simple, Efficient, and Production Ready Training Library

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has emerged as a prominent paradigm for training large language models and multimodal systems. Despite the notable advances enabled by existing RLHF training frameworks, significant challenges remain to scale to complex multimodal workflows and adapt to dynamic workloads. In particular, current systems often encounter limitations related to controller scalability when managing large models, as well as inefficiencies in orchestrating intricate RLHF pipelines, especially in scenarios that require dynamic sampling and resource allocation. In this paper, we introduce WeChat-YATT (Yet Another Transformer Trainer in WeChat), a simple, scalable, and balanced RLHF training framework specifically designed to address these challenges. WeChat-YATT features a parallel controller programming model that enables flexible and efficient orchestration of complex RLHF workflows, effectively mitigating bottlenecks associated with centralized controller architectures and facilitating scalability in large-scale data scenarios. In addition, we propose a dynamic placement schema that adaptively partitions computational resources and schedules workloads, thereby significantly reducing hardware idle time and improving GPU utilization under variable training conditions. We evaluate WeChat-YATT across diverse experimental scenarios, demonstrating its substantial throughput improvements over state-of-the-art RLHF training frameworks. Furthermore, WeChat-YATT has been successfully deployed to train models that support WeChat product features for a large-scale user base, underscoring its effectiveness and robustness in real-world applications. We have made WeChat-YATT publicly available at https://www.github.com/tencent/WeChat-YATT.
@@ 1 Introduction
Transformer-based large models and their multimodal extensions have achieved remarkable progress, demonstrating state-of-the-art performance across a wide range of natural language processing, understanding, and generation tasks [1, 2, 3, 4, 5, 6, 7, 8, 9]. These models, characterized by their massive parameter counts and ability to process extremely long context sequences, have become the foundation for many advanced AI applications. However, the unprecedented scale of LLMs has introduced significant challenges for efficient training, and traditional data-parallel training strategies are often insufficient to handle the memory and computational demands of such large models. To address these challenges, the research community has developed a variety of hybrid parallelism techniques [10, 11, 12, 13, 14]. These approaches enable the training of models that would otherwise exceed the memory and compute capabilities of individual devices, by partitioning both the model and the data across multiple GPUs or nodes.
@@
The integration of Reinforcement Learning from Human Feedback (RLHF) has significantly transformed the training paradigms of large language models [15, 16, 17]. By embedding human preferences into the optimization process, RLHF requires the coordination of multiple models—such as reward models, policy models, and reference models—within a unified training pipeline [18, 19, 20]. This multi-model architecture substantially amplifies the number of parameters and the volume of intermediate activations that must be handled, thereby straining the capabilities of current distributed training infrastructures. The scalability challenges are further exacerbated in multimodal RLHF scenarios. The incorporation of diverse data modalities (e.g., text, images, audio) not only increases the heterogeneity and volume of training data but also intensifies the demands on inter-node communication and memory bandwidth. As a result, the system faces pronounced communication overheads and memory bottlenecks, hindering its ability to scale efficiently to larger datasets and model sizes. Addressing these bottlenecks is critical for facilitating the practical deployment of RLHF in increasingly complex and data-rich environments.
@@
Sample efficiency is a critical factor in the RLHF training process, as it directly influences convergence rates and overall model performance [21]. While flexible sampling strategies have been shown to accelerate convergence, most existing RLHF training frameworks lack robust support for efficient and adaptive sampling mechanisms. Although hybrid parallelism theoretically enables substantial scalability, practical implementations often encounter significant communication overhead as the degree of parallelism increases. Specifically, as the number of models and the length of context sequences expand, the costs associated with synchronizing parameters and activations across distributed devices can escalate rapidly, thereby diminishing the anticipated benefits of parallelization. To address these challenges, recent work such as Shen et al. [22] has conceptualized the RLHF training workflow as a control flow problem, employing techniques such as multi-cluster, multi-model partial colocated, and memory swapping to alleviate peak GPU memory consumption. Similarly, other systems [23, 24] have explored alternative strategies, with Sheng et al. [24] further enhancing throughput by integrating both single- and multi-controller architectures alongside statically optimized placement schemes. These advancements underscore the ongoing need for innovative approaches to improve sampling efficiency and resource utilization in large-scale RLHF training.
@@
The proposed WeChat-YATT advances the state-of-the-art in RLHF training by effectively addressing the critical bottlenecks of controller scalability and resource placement. We hope that WeChat-YATT will serve as a robust foundation for both academic research and industrial applications, fostering further innovation and collaboration within the community.
@@ 2.1 Distributed Training of Deep Models
Data Parallelism involves splitting the training data across multiple devices (such as GPUs), where each device holds a copy of the model and processes a different mini-batch of data. Gradients are averaged (or summed) across devices to update the model parameters synchronously [25, 26]. Tensor parallelism splits individual layers or tensors of the model across multiple devices. Each device is responsible for computing a portion of the operations within a layer, enabling the training of very large models that cannot fit into a single device’s memory [10]. Pipeline parallelism divides the model into sequential stages, with each stage placed on a different device. Mini-batches are divided into micro-batches and passed through the pipeline, allowing different devices to work on different parts of the input simultaneously, improving hardware utilization [11, 27, 28].
ZeRO is an optimization technique that partitions model states (such as optimizer states, gradients, and parameters) across devices to minimize memory redundancy. This enables the training of much larger models by reducing the memory consumption per device [29, 30].
@@
Expert Parallelism is used in Mixture-of-Experts (MoE) models, where different subsets of the model (experts) are distributed across devices. During training and inference, only a subset of experts is activated for each input, allowing efficient scaling and utilization of resources [13, 33].
@@ 2.2 RLHF Workflow
Most RLHF workflows largely follow this 4-stage workflow [18, 19, 40, 16, 15, 41, 20](Sometimes, the first three stages are referred to as ’rollout’ (Figure 1(a)):
@@
    Stage 3(Preparation): Using generated responses and rewards, the critic computes their values, the policy and reference policy compute their reference log probabilities.
@@
To simplify implementation, most systems use a single controller to manage data transfer, intermediate results, and workflow orchestration through a central point [23, 24]. However, this centralized approach can sometimes become a bottleneck, especially when the trainer itself has high memory requirements.
@@ 2.3 RLHF Placement
  • Co-locate: Multiple models are placed on the same set of GPUs, sharing GPU memory. To avoid out-of-memory (OOM) errors, which can easily occur if large models are run simultaneously, these colocated models are executed one after another in a time-sharing fashion.
Although Sheng et al. [24] proposed methods for calculating placement based on workload, both the open source implementations of OpenRLHF [23] and VeRL [24] adopt a co-locate all models strategy. The partial colocated approach introduces overhead from swapping among multiple models. For example, swapping a 32B model could take nearly a minute, and updating weights may take tens of seconds. In typical GRPO training, this overhead is not a major issue, since the primary bottlenecks in RLHF training are the generation and the forward/backward passes of modeling. Compared to the tens of minutes spent on rollout and training, model swapping is not the system bottleneck.
Some works [44, 45] have adopted a coexisting approach, which effectively avoids the overhead of model swapping. However, to mask device idle time during workflow state transitions, these methods introduce varying degrees of asynchronous overlap between the rollout and training stages. This can lead to suboptimal model convergence; for example, Fu et al. [44] introduces staleness-aware training to address this issue. In our experience, many algorithm engineers find that this trade-off adds extra cognitive burdens.
@@ 3.1 Overview
Figure 2 presents a high-level overview of the proposed WeChat-YATT framework. WeChat-YATT supports two distinct resource placement architectures: partial colocated and full colocated.
In the partial colocated architecture, the Sampler and the GenRM (Generative Reward Model) components are deployed independently and interact asynchronously. The allocation of computational resources to these two components directly influences throughput during the rollout phase, a placement method that will be examined in detail in Section 3.3. Within this architecture, the actor model leverages all available machine resources across the cluster, while parallel controllers correspond to the ranks of model-parallel worker heads associated with the Actor. The training loop is initiated by wake-up remote procedure calls (RPCs), activating both the Sampler and GenRM. Subsequently, parallel controllers retrieve data from the distributed GDataset, thereby mitigating excessive memory consumption on any single machine. Notably, both the Sampler and GenRM operate asynchronously: upon completing each microbatch, the Sampler immediately transmits the generated data to the GenRM without impeding ongoing generation processes. Once the rollout phase concludes, the Actor instructs other components to enter a sleep state, offloads their memory, and subsequently loads the actor model to initiate the training phase.
In contrast, the full colocated architecture adopts a more integrated approach, wherein all components share the resources of all machines. Under this configuration, the GenRM must await the Sampler’s completion of the entire sample generation process. Following this, the Sampler offloads its parameters to disk, and the GenRM is loaded into memory to initiate the reward computation process. Upon completing reward computation, the Actor is loaded into memory, while the GenRM is offloaded. This workflow is inherently synchronous, with transitions between components occurring only after the preceding process is fully completed. Consequently, in scenarios characterized by frequent long-tail effects, the full colocated architecture may not offer optimal performance.
@@ 3.3 Dynamic Placement
However, dynamic sampling scenarios introduce additional complexity. As described in Yu et al. [40], prompts with accuracy scores of 1 or 0 (i.e., batches with uniformly high or low rewards) are filtered out, triggering resampling. This process necessitates additional rounds of generation and reward computation, as illustrated in Figure 1(b).
@@
To address the challenges associated with model swapping overhead and the exacerbated long-tail effect, we develope a dynamic placement strategy that integrates dynamical resource allocation in the training process under patial colocated WeChat-YATT paradigm.
@@
When the workload is dynamic, static estimation methods such as those proposed in Sheng et al. [24] become inadequate for determining optimal resource placement. It is inherently difficult to predict how the output length of the model will evolve over the course of training. Moreover, as the model’s reasoning responses become longer, it becomes increasingly challenging to anticipate the corresponding generation length required for reward computation. This unpredictability further complicates the design of efficient placement strategies in dynamic training environments.
@@ 4.4 Workload Balancing
As noted in Wang et al. [54] and Jiang et al. [55], workload imbalance can create significant bubbles during training, becoming a major bottleneck. When sequence lengths are long, the main training bottleneck is attention computation, whose complexity is $s^{2}$ for a sequence of length $s$, while the cost of packing is much lower. For example, if a sequence is formed by concatenating two sequences of length $\frac{s}{2}$, the total computation is $\frac{s^{2}}{2}$, leading to severe load imbalance and device underutilization. This issue is even more pronounced with post-training data, which often varies greatly in length.
Compared to complex combinatorial optimization approaches, we find a much simpler solution for workload balancing: instead of sequence packing, we simply sort data by simulated workload ($s^{2}$), greatly simplifying the process and even improving computational efficiency.
@@ 4.5 Distributed Attention
Currently, most implementations achieve context parallelism using ring attention for both causal and full attention. To support more complex attention masks (e.g., Gemma-3 [56]), inspired by Llama-3 [2], we implement distributed attention based on CCL all-gather, where we first all-gather the key (K) and value (V) tensors, and then compute attention output for the local query (Q) tensor chunk.
@@ 5.1.1 Co-locate Generative Reward Model
In the full colocated WeChat-YATT setup, we conduct experiments on the generative reward model compared with VeRL under two distinct scenarios. The first scenario features a balanced parameter size between the generative reward model and the actor model, while the second scenario considers a highly unbalanced configuration. As VeRL currently supports only external calls to the generative reward model, we allocate computational resources in proportion to the parameter size ratio between the generative reward and actor models.
For the first comparation scenario setup, we select the VeRL Generative Reward demo. In this configurationof VeRL, GenRM-CI-Test-1.5B serves as the external reward model, utilizing 8 GPUs for inference. SGlang is employed as the inference engine, with a data parallelism size (dp-size) of 8. An additional 8 GPUs are allocated to deploy the actor model, Qwen2.5-3B-Instruct, with a tensor parallelism size (tp-size) of 2. For the training backend, we adopt the state-of-the-art Megatron-LM framework. To ensure a fair comparison, the WeChat-YATT setup mirrors the VeRL configuration, employing a full colocated deployment with 16 GPUs.
@@
We further evaluate performance in a scenario involving a large generative reward model, Qwen2.5-Math-72B-Instruct, paired with a smaller actor model, Qwen2.5-Math-1.5B. This configuration corresponds to the generative reward model demonstration provided in our open-source codebase. For the VeRL setup, 24 GPUs are allocated for inference, while an additional 8 GPUs are dedicated to actor model rollout generation and training. In contrast, WeChat-YATT employ a full colocated configuration utilizing 32 GPUs, leveraging an offload/onload strategy. Megatron-LM is used as the training backend in both cases.
@@ 5.1.2 Dynamic Sampling
To assess the performance of the partial colocated WeChat-YATT framework in a dynamic sampling context, we conduct experiments in which varying sample rejection rates are applied during the rollout phase. Consistent with previous experimental settings, GenRM-CI-Test-1.5B is employed as the generative reward model, while Qwen2.5-3B-Instruct served as the actor model. The sample rejection rates are varied from 10% to 40%, and the framework need to re-sample rejected rollouts until all samples are accepted. The primary metric of interest is the time consumption observed during the rollout phase under these different rejection rates.
As shown in Table LABEL:tab:reject_rates, the rollout time for both frameworks increases as the sample rejection rate rises. However, the partial colocated WeChat-YATT consistently demonstrates lower rollout time across all rejection rates when compared to the full colocated approach. Specifically, the partial colocated framework achieves a reduction in rollout time usage ranging from 16.2% to 24.1% relative to the full colocated configuration.
This performance improvement can be attributed to the asynchronous interaction between the actor rollouts and the generative reward model in the partial colocated design. In this setup, each completed microbatch of actor rollouts is immediately dispatched to the standalone generative reward model in an asynchronous manner. This process effectively mitigates the long-tail latency that often constrains throughput in the full colocated WeChat-YATT framework. Furthermore, the frequent context switching between actor rollouts and the generative model in the full colocated setup introduces additional overhead, which is substantially reduced in the partial colocated architecture.